# Gravity from entropy

### v5

@@ I Introduction
In this work, a continuum modified theory of gravity based on a statistical mechanics action is considered. This theory treats the metric at each point of spacetime as a renomalizable density matrix, or more precisely a local quantum operator. This central idea is relating geometry with the mathematical foundation of quantum field theory [33] and is inspired by the used of von Neumann algebras in explaining entanglement in field theories [17, 18] and quantum gravity  [34, 35, 36]. While the geometry of spacetime is defined through its associated metric, the interplay between matter fields and geometry is captured by the metric induced by the matter fields which describes how the matter fields effectively curves spacetime. Embracing a statistical mechanics approach to gravity, this work interprets these metric tensors as quantum operators and postulates an action for gravity given by the quantum relative entropy between the metric of the manifold and the metric induced by the matter fields. From the mathematical point of view, the quantum relative entropy proposed in this work is strictly related to the Araki quantum relative entropy for von Neumann algebras [37, 38, 17, 39]. From the physics point of view, the proposed action fully describes how matter curves geometry and how geometry affects the matter fields.
@@
This work greatly expands on previous results [44] obtained in the discrete setting by the same author. On one side, here a continuum and fully Lorentz invariant theory is proposed. This progress is based on the development of the suitable mathematical framework to define the Lorentz invariant entropy and cross-entropy between the metric of the spacetime and the metric induced by the matter fields. On the other side, here the relation of this statistical mechanics/information theory action with the Einstein-Hilbert action [46, 47] is established defining a clear connection to gravity. Two fundamental aspects of this work that are not present in Ref. [44], are relevant to clearly relate this approach to gravity. First, this work considers a local theory, defining the entropy of the metric at each point of spacetime, while the previous work considers the entropy associated to the full metric of the higher-order network. The present local theory allows a closer connection to gravity and constitutes a step forward to establish the connection between this approach and the quantum theory of entanglement [17]. Secondly, by adopting the continuum limit, in this work the intrinsic difficulty related to the definition of the curvature of networks, simplicial and cell complexes is avoided.
@@
This work is structured as follows. In Sec. II we provide the motivation of the proposed theory and we discuss preliminary results on an instructive warm-up scenario. In Sec. III we outline the proposed theoretical framework, we postulate the entropic action for gravity and we derive the corresponding modified Einstein equations. In Sec.IV we provide the concluding remarks. The paper also includes three Appendices discussing possible extensions of the proposed theoretical framework, and providing the mathematical background and all the necessary details regarding the notation used Sec. III .
@@ II.2 Warm-up scenario: Entropic action
Note that the problem of defining eigenvalues and entropy of tensors is a topic of intense research and similar definitions are been provided for instance in the theory of elasticity [50] and in applied tensor analysis [51] (although not defined in a Lorentz invariant way). In the present warm-up scenario we will define the entropy and the quantum relation entropy of different metrics taking the form of rank-2 tensors. As a reference for the notation used, the reader can refer to Table 1. Recalling that the metric $g$ has all the eigenvalues equal to one, it follows that the entropy of of the metric $g$ is null, i.e.
@@
$\mathcal{L}\equiv-\mbox{Tr}_{M}\ln\ {\bf G}{g}^{-1}=-\sum_{\mu}\ln(\mu),$ (12)
where with $\mbox{Tr}_{M}\ln\ {\bf G}{g}^{-1}$ we indicate the usual trace of the logarithm of the matrix $\ln\ {\bf G}{g}^{-1}$ and with $\mu$ the eigenvalues of the $2$-tensor ${\bf G}$ as defined in Eq.(2).
@@
Although we are not aware of previous interpretations of the metrics as quantum operators, it is well known that the quantum relative entropy can also be defined among quantum operators that generalize density matrices. In particular the Araki quantum relative entropy [37, 38]. is defined between quantum operators, that like our metric matrices admit a finite trace at each point of the manifold but this trace might be not unitary and be a function of the considered point of the manifold.
While we leave the discussion of the relation of our entropy action to the Araki quantum relative entropy to Appendix C, a series of physical observations are here in place to justify our interpretation of the metrics matrices as quantum operators. As remarked previously we treat the metric tensors as quantum operators or effective density matrices. The main differences between the metric tensor and the density operators include the fact that we require that the metric matrices are invertible and we do not require that they have unitary trace. The requirement of treating exclusively invertible metrics $g$ and ${\bf G}$ is dictated by our desire to define the entropy in a Lorentz invariant way. In fact here we desire to treat the metrics $g$ and ${\bf G}$ on the same footing of their inverse $g^{-1}$ and ${\bf G}^{-1}$. It might be argued that also our relaxation of the requirement to have density matrices of trace one, has similar roots. In fact if we require $g$ and ${\bf G}$ to have unitary trace at each point of the manifold their inverse in general will not have a unitary trace.
@@ II.3 Warm-up scenario: scalar matter fields
Adopting this notation we observe that the logarithm of the induced metric $\ln{\bf G}$ and the logarithm of $\ln{\bf G}^{-1}$ are given by
$\ln{\bf G}$ $=$ $f(|{\bm{\nabla}}\phi|^{2}){\mathcal{G}},$
$\ln{\bf G}^{-1}$ $=$ $-f(|{\bm{\nabla}}\phi|^{2}){\mathcal{G}}^{-1},$ (18)
@@ III The entropic theory of matter fields coupled to geometry
$\tilde{\bf G}=G_{(0)}+[G_{(1)}]_{\mu\nu}dx^{\mu}\otimes dx^{\nu}\oplus[G_{[2]}]_{\mu\nu\rho\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma})$ covariant topological metric induced by the matter fields
@@ III.1 Topological matter fields and their associated metrics
$\bra{\Phi}=\bar{\phi}\oplus\bar{\omega}_{\mu}dx^{\mu}+\bar{\zeta}_{\mu\nu}dx^{\mu}\wedge dx^{\nu}.$ (25)
@@
$\tilde{\bf G}$ $=$ $G_{(0)}+[G_{(1)}]_{\mu\nu}dx^{\mu}\otimes dx^{\nu}$ (34)
$\oplus[G_{[2]}]_{\mu\nu\rho\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma}).$
Proceeding as in the previous case we might wish to define the metric induced by the topological field as
@@
In our interpretation of the metric as a quantum density matrix, this would correspond to density matrix corresponding to a pure state $D\ket{\Phi}$. However, if the induced metric $\tilde{\bf G}$ is interpreted as a density matrix, it is natural to add further terms in the metric induced by the topological matter field $\tilde{\bf G}$. These terms will depend on the topological field and the geometry of the space which will allow us to describe mixed states. First, we introduce in $\tilde{\bf G}$ a term proportional to |$\ket{\Phi}\bra{\Phi}$ |. Specifically we introduce a term $(m^{2}+\xi R)\ket{\Phi}\bra{\Phi}$ | where $R$ is the Ricci scalar and $\xi$ might include the case of conformal coupling $\xi=1/6$. This term can be also interpreted as a projector. Secondly, we introduce a term depending explicitly on the curvature of the manifold. Since $\tilde{\bf G}$ involves the metric for scalars, vectors and bivectors on equal footing, it is natural to consider a further term involving $\tilde{\bm{\mathcal{R}}}$ given by the direct sum of the Ricci scalar $R$, the Ricci tensor of elements $R_{\mu\nu}$ and the Riemann tensor of elements $R_{\mu\nu\rho\sigma}$, i.e.
@@
From these considerations, it follows that in the induced metric $\tilde{\bf G}$, we will substitute the term $D\ket{\Phi}\bra{\Psi}D$ with $\tilde{\bf M}$ defined as
@@ III.2 Entropic topological and geometrical action
(see Figure 1 for an illustration of the physical model beyond the choice of this Lagrangian). Since we have $H=0$ we obtain
@@
$\bra{\Phi}{D}\tilde{g}^{-1}{D}\ket{\Phi}$ $=$ $|\nabla\phi|^{2}+|\nabla^{\mu}\omega_{\mu}|^{2}+\Big{|}\nabla^{\sigma}\bar{\omega}_{\sigma\mu}\Big{|}^{2}$ (47)
@@
In the interesting limit in which $\omega_{\mu}=0$ and $\zeta_{\mu\nu}=0$ and the bosonic matter field is only scalar, Eq.(46) reduces to the widely studied Einstein-Hilbert action with zero cosmological constant coupled with the scalar field [53],
@@ III.3 Introduction of G-field $\tilde{\mathcal{G}}$ and the field $\tilde{\Theta}$
This Lagrangian is now linear in $\tilde{\bf G}\tilde{g}^{-1}$. Note that a conservative interpretation of this transformation will not give a physical interpretation to the auxiliary fields. Thus, according to this point of view, the auxiliary fields are not changing the physics of the original Lagrangian and this theory might be prone to Ostrogradsky instability. However, we known from statistical physics that Lagrangian multipliers can also acquire a rather physical meaning. For instance the temperature and the chemical potential are Lagrangian multipliers that are measurable. Here we would like to embrace this point of view and give a physical meaning to the introduced Lagrangian multipliers as physical fields defined on our manifold $\mathcal{K}$. Interpreting in this way the Lagrangian multipliers changes the physics of the problem as the phase space associated to the equations of motion acquire new dimensions and the fixed points as well as the initial value problem will involve the newly introduced fields. In this latter interpretation, introducing the auxiliary fields and giving a physical meaning to them as encoding for new fields, we have reduced our theory to a theory linear in $\tilde{\bm{\mathcal{R}}}$. In particular, as we will see below, the equations of motion will involve at most second order derivatives of the fields, thus the model will avoid the Ostrogradsky instability.
@@ III.5 Modified gravity
We now turn to the equations for modified gravity considering variation of our action $\tilde{\mathcal{S}}$ with respect to the $g$ and the fields $\tilde{\bm{\Theta}}$ and $\tilde{\bm{\mathcal{G}}}$.
@@ III.5.2 Variation with respect to $\tilde{\bm{\Theta}}$
$\Lambda_{\mathcal{G}}=\frac{1}{2\beta}\Big{(}\tilde{\bm{\mathcal{G}}}-\tilde{\bf I}-\mbox{Tr}_{F}\ln\tilde{\bm{\mathcal{G}}}\Big{)}.$ (65)
@@ III.5.3 Variation with respect to ${g}$
The modified Einstein equations are obtained by performing the variation of the action $\hat{\mathcal{S}}$ with respect to $g$. Let us define the stress-energy tensor $\bm{\mathcal{T}}$ in the usual way with elements ${\mathcal{T}}_{\mu\nu}$ given by
@@
${R}^{\mathcal{G}}_{\mu\nu}$ $=$ ${\mathcal{G}_{(0)}}R_{\mu\nu}+{[{\mathcal{G}_{(1)}}]}_{\mu}^{\ \rho}R_{\rho\nu}$ (68)
$+2{[{\mathcal{G}_{(2)}}]}_{\mu\eta}^{\ \ \rho_{1}\rho_{2}}g^{\sigma\eta}R_{\rho_{1}\rho_{2}\nu\sigma},$
@@
${{\mathcal{D}}}_{\mu\nu}$ $=$ $(\nabla^{\rho}\nabla_{\rho}g_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}){\mathcal{G}_{(0)}}+2\nabla^{\eta}\nabla^{\rho}{[\mathcal{G}_{(2)}]}_{\mu\rho\nu\eta}$ (69)
@@
$-\nabla^{\rho}\nabla_{\nu}{[\mathcal{G}_{(1)}]}_{(\rho\mu)}.$
@@ IV Conclusions
This work was partially supported by a grant from the Simons Foundation. The author would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Hypergraphs: Theory and Applications, where work on this paper was undertaken. This work was supported by EPSRC grant EP/R014604/1.

### v7

@@ I Introduction
In this work, a continuum modified theory of gravity based on a statistical mechanics action is considered. This theory treats the metric at each point of spacetime as a “renomalizable” density matrix, or more precisely a local quantum operator. This central idea is relating geometry with the mathematical foundation of quantum field theory [33] and is inspired by the used of von Neumann algebras in explaining entanglement in field theories [17, 18] and quantum gravity  [34, 35, 36]. While the geometry of spacetime is defined through its associated metric, the interplay between matter fields and geometry is captured by the metric induced by the matter fields which describes how the matter fields effectively curve spacetime. Embracing a statistical mechanics approach to gravity, this work interprets these metric tensors as quantum operators and postulates an action for gravity given by the quantum relative entropy between the metric of the manifold and the metric induced by the matter fields. From the mathematical point of view, the quantum relative entropy proposed in this work is strictly related to the Araki quantum relative entropy for von Neumann algebras [37, 38, 17, 39]. From the physics point of view, the proposed action fully describes how matter curves geometry and how geometry affects the matter fields.
@@
This work greatly expands on previous results [44] obtained in the discrete setting by the same author. On one side, here a continuum and fully Lorentz invariant theory is proposed. This progress is based on the development of the suitable mathematical framework to define the Lorentz invariant entropy and cross-entropy between the metric of the spacetime and the metric induced by the matter fields. On the other side, here the relation of this statistical mechanics/information theory action with the Einstein-Hilbert action [46, 47] is established defining a clear connection to gravity. Two fundamental aspects of this work that are not present in Ref. [44], are relevant to clearly relate this approach to gravity. First, this work considers a local theory, defining the entropy of the metric at each point of spacetime, while the previous work considers only the entropy associated to the full metric of the higher-order network. The present local theory allows a closer connection to gravity and constitutes a step forward to establish the connection between this approach and the quantum theory of entanglement [17]. Secondly, by adopting the continuum limit, in this work the intrinsic difficulty related to the definition of the curvature of networks, simplicial and cell complexes is avoided.
@@
This work is structured as follows. In Sec. II we provide the motivation of the proposed theory and we discuss preliminary results on an instructive warm-up scenario. In Sec. III we outline the proposed theoretical framework, we postulate the entropic action for gravity and we derive the corresponding modified Einstein equations. In Sec.IV we provide the concluding remarks. The paper also includes three Appendices discussing possible extensions of the proposed theoretical framework, providing the mathematical background and all the necessary details regarding the notation used Sec. III, and establishing the connection of the present theory with the theory of local quantum operators and the Araki entropy.
@@ II.2 Warm-up scenario: Entropic action
Note that the problem of defining eigenvalues and entropy of tensors is a topic of intense research and similar definitions have been provided for instance in the theory of elasticity [50] and in applied tensor analysis [51] (although not defined in a Lorentz invariant way). In the present warm-up scenario we will define the entropy and the quantum relation entropy of different metrics taking the form of rank-2 tensors. As a reference for the notation used, the reader can refer to Table 1. Recalling that the metric $g$ has all the eigenvalues equal to one, it follows that the entropy of the metric $g$ is null, i.e.
@@
$\mathcal{L}\equiv-\mbox{Tr}_{M}\ln\ {\bf G}{g}^{-1}=-\sum_{\lambda^{\prime}}\ln(\lambda^{\prime}),$ (12)
where with $\mbox{Tr}_{M}\ln\ {\bf G}{g}^{-1}$ we indicate the usual trace of the logarithm of the matrix $\ln\ {\bf G}{g}^{-1}$ and with $\lambda^{\prime}$ the eigenvalues of the $2$-tensor ${\bf G}$ as defined in Eq.(2).
@@
Although we are not aware of previous interpretations of the metrics as quantum operators, it is well known that the quantum relative entropy can also be defined among quantum operators that generalize density matrices. In particular, in the fundamental theory of quantum operators algebras [37, 38], quantum operators generalize the notion of density matrices in an analogous way of the metric tensors adopted here. In particular in this theory, quantum operators, like our metric matrices, might admit a finite trace at each point of the manifold but this trace might be not unitary and might be a function of the considered point of the manifold. For these quantum operators the Araki quantum relative entropy [37, 38] is defined. This entropy reduces to the von Neumann entropy when the quantum operators reduces to a density matrices of unitary trace but is defined also among quantum operators of non unitary trace.
While we leave the discussion of the relation of our entropic action to the Araki quantum relative entropy to Appendix C, a series of physical observations are here in place to justify our interpretation of the metrics matrices as quantum operators. As remarked previously we treat the metric tensors as quantum operators or effective density matrices. The main differences between the metric tensor and the density operators include the fact that we require that the metric matrices are invertible and we do not require that they have unitary trace at each point of the manifold. The requirement of treating exclusively invertible metrics $g$ and ${\bf G}$ is dictated by our desire to define the entropy in a Lorentz invariant way. In fact here we desire to treat the metrics $g$ and ${\bf G}$ on the same footing of their inverse $g^{-1}$ and ${\bf G}^{-1}$. It might be argued that also our relaxation of the requirement to have density matrices of trace one, has similar roots. In fact if we require $g$ and ${\bf G}$ to have unitary trace at each point of the manifold their inverse as well as their dual (defined in Appendix C) in general will not have a unitary trace.
@@ II.3 Warm-up scenario: scalar matter fields
Adopting this notation we observe that the logarithm of the induced metric $\ln{\bf G}$ and the logarithm of $\ln{\bf G}^{-1}$ have elements
$\left[\ln{\bf G}\right]_{\mu\nu}$ $=$ $f(|{\bm{\nabla}}\phi|^{2}){M}_{\mu\nu},$
$\left[\ln{\bf G}^{-1}\right]^{\mu\nu}$ $=$ $-f(|{\bm{\nabla}}\phi|^{2}){M}^{\mu\nu},$ (18)
@@ III The entropic theory of matter fields coupled to geometry
$\tilde{\bf G}=G_{(0)}\oplus[G_{(1)}]_{\mu\nu}dx^{\mu}\otimes dx^{\nu}\oplus[G_{(2)}]_{\mu\nu\rho\sigma}(dx^{\mu}\wedge dx^{\nu})\otimes(dx^{\rho}\wedge dx^{\sigma})$ Covariant topological metric induced by the matter fields
@@ III.1 Topological matter fields and their associated metrics
$\bra{\Phi}=\bar{\phi}\oplus\bar{\omega}_{\mu}dx^{\mu}\oplus\bar{\zeta}_{\mu\nu}dx^{\mu}\wedge dx^{\nu}.$ (25)
@@
$\tilde{\bf G}$ $=$ $G_{(0)}\oplus[G_{(1)}]_{\mu\nu}dx^{\mu}\otimes dx^{\nu}$ (33)
Proceeding as in the warm-up scenario we might wish to define the metric induced by the topological field as
@@
In our interpretation of the metric as a quantum density matrix, this would correspond to density matrix corresponding to a pure state $D\ket{\Phi}$. However, if the induced metric $\tilde{\bf G}$ is interpreted as a density matrix, it is natural to add further terms in the metric $\tilde{\bf G}$. These terms will depend on the topological field and the geometry of the space which will allow us to describe mixed states. First, we introduce in $\tilde{\bf G}$ a term proportional to |$\ket{\Phi}\bra{\Phi}$ |. Specifically we introduce a term $(m^{2}+\xi R)\ket{\Phi}\bra{\Phi}$ | where $R$ is the Ricci scalar and $\xi$ might include the case of conformal coupling $\xi=(d-2)/(4(d-1))$. This term can be also interpreted as a projector. Secondly, we introduce a term depending explicitly on the curvature of the manifold. Since $\tilde{\bf G}$ involves the metric for scalars, vectors and bivectors on equal footing, it is natural to consider a further term involving $\tilde{\bm{\mathcal{R}}}$ given by the direct sum of the Ricci scalar $R$, the Ricci tensor of elements $R_{\mu\nu}$ and the Riemann tensor of elements $R_{\mu\nu\rho\sigma}$, i.e.
@@
From these considerations, it follows that in the induced metric $\tilde{\bf G}$, we will substitute the term $D\ket{\Phi}\bra{\Phi}D$ with $\tilde{\bf M}$ defined as
@@ III.2 Entropic topological and geometrical action
(see Figure 1 for an illustration of the physical model beyond the choice of this Lagrangian). Since we have $\tilde{H}=0$ we obtain
@@
$\bra{\Phi}{D}\tilde{g}^{-1}{D}\ket{\Phi}$ $=$ $|\nabla\phi|^{2}+|\nabla^{\mu}\omega_{\mu}|^{2}+\Big{|}\nabla^{\rho}{\zeta}_{\rho\mu}\Big{|}^{2}$ (46)
@@
In the interesting limit in which $\omega_{\mu}=0$ and $\zeta_{\mu\nu}=0$ and the topological matter field is only scalar, Eq.(45) reduces to the widely studied Einstein-Hilbert action with zero cosmological constant coupled with the scalar field [53],
@@ III.3 Introduction of G-field $\tilde{\mathcal{G}}$ and the field $\tilde{\Theta}$
This Lagrangian is now linear in $\tilde{\bf G}\tilde{g}^{-1}$. Note that a conservative interpretation of this transformation will not give a physical interpretation to the auxiliary fields. Thus, according to this point of view, the auxiliary fields are not changing the physics of the original Lagrangian and this theory might be prone to Ostrogradsky instability. However, we known from statistical physics that Lagrangian multipliers can also acquire a rather physical meaning. For instance the temperature and the chemical potential are Lagrangian multipliers that are measurable. Here we would like to embrace this point of view and give a physical meaning to the introduced Lagrangian multipliers as physical fields defined on our manifold $\mathcal{K}$. Interpreting in this way the Lagrangian multipliers changes the physics of the problem as the phase space associated to the equations of motion acquire new dimensions and the fixed points as well as the initial value problem will involve the newly introduced fields. In this latter interpretation, introducing the auxiliary fields and giving a physical meaning to them as encoding for new fields, we have reduced our theory to a theory linear in $\tilde{\bm{\mathcal{R}}}$. In particular, as we will see below, the equations of motion will involve at most second order derivatives of the fields, thus the model might avoid the Ostrogradsky instability. A definitive answer to the stability question would follow from further analysis (for example a Hamiltonian analysis).
@@ III.5 Modified gravity
We now turn to the equations for modified gravity considering variation of our action $\tilde{\mathcal{S}}$ with respect to the metric $g$ and to the fields $\tilde{\bm{\Theta}}$ and $\tilde{\bm{\mathcal{G}}}$.
@@ III.5.2 Variation with respect to $\tilde{\bm{\Theta}}$
$\Lambda_{\mathcal{G}}=\frac{1}{2\beta}\mbox{Tr}_{F}\Big{(}\tilde{\bm{\mathcal{G}}}-\tilde{\bf I}-\ln\tilde{\bm{\mathcal{G}}}\Big{)}.$ (64)
@@ III.5.3 Variation with respect to ${g}$
The modified Einstein equations are obtained by performing the variation of the action $\tilde{\mathcal{S}}$ with respect to $g$. Let us define the stress-energy tensor $\bm{\mathcal{T}}$ in the usual way with elements ${\mathcal{T}}_{\mu\nu}$ given by
@@
${R}^{\mathcal{G}}_{\mu\nu}$ $=$ ${\mathcal{G}_{(0)}}R_{\mu\nu}+{[{\mathcal{G}_{(1)}}]}_{\mu}^{\ \rho}R_{\rho\nu}-{[\mathcal{G}_{(2)}]}_{\rho_{1}\rho_{2}\mu\eta}R_{\nu}^{\ \eta\rho_{1}\rho_{2}}$ (67)
$+2{[{\mathcal{G}_{(2)}}]}_{\mu}^{\ \eta\rho_{1}\rho_{2}}R_{\rho_{1}\rho_{2}\nu\eta},$
@@
${{\mathcal{D}}}_{\mu\nu}$ $=$ $(\nabla^{\rho}\nabla_{\rho}g_{\mu\nu}-\nabla_{\mu}\nabla_{\nu}){\mathcal{G}_{(0)}}-\nabla^{\rho}\nabla_{\nu}{[\mathcal{G}_{(1)}]}_{(\rho\mu)}$ (68)
@@
$+\nabla^{\eta}\nabla^{\rho}{[\mathcal{G}_{(2)}]}_{\mu\rho\nu\eta}+\nabla^{\rho}\nabla^{\eta}{[\mathcal{G}_{(2)}]}_{\eta\mu\rho\nu}$
$+\frac{1}{2}[\nabla^{\rho},\nabla^{\eta}]{[\mathcal{G}_{(2)}]}_{\rho\eta\mu\nu}.$
@@ IV Conclusions
This work was partially supported by a grant from the Simons Foundation. The author would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Hypergraphs: Theory and Applications, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1.